A high-end watch company records fixed costs of $175,000 to make gold watches. If the total...

  1. A high-end watch company records fixed costs of $175,000 to make gold watches. If the total cost to make 20 watches is $560,000, what is the marginal cost of a 21st gold watch, rounded to the nearest whole dollar?
    1. $18,333
    2. $19,250
    3. $194,250
    4. $385,000
    5. Some other amount
Answer #1

Ans: The answer of the above mentioned question is option b. $19,250.

Explanation: In simple words, Marginal cost is the additional cost that is to be incurred to produce one additional unit I.e. the change in total cost by producing one additional unit.

Formula: The formula for Marginal costs is as follows:

Marginal cost = change in total cost/change in quantity.

In the given question, high-end watch company produces 20 watches whose fixed cost as given is $175,000 and the total cost is $560,000. Hence, the variable production cost to produce 20 watched is $385,000 (I.e. Total cost- Fixed cost= 560000-175000).

Therefore, the variable cost per watch is Total Variable cost/no. Of watched.

I.e. 385000/20 = $19,250.

Since, the fixed costs remains same the additional cost of producing one additional unit will be $19,250.

Alternative through formula: (For 21 watches)

Fixed cost: $175,000

Variable cost: $404,250 (I.e. $19,250 * 21)

Total Cost= $579,250.

Therefore, Marginal Cost = Change in total cost/change in quantity.

I.e. (579250-560000)/ (21-20) = $19,250.

Other 3 options are not correct since they are not calculated as per Marginal Cost definition. Thankyou.